Operations Report – February 18th

Crew 205 Operations Report 18 – FEB – 2019

SOL: 8

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1535 (18thFeb), turned off at 1100 (18thFeb)

Solar— SOC % – Turned off Gen (18th Feb) 95%, Turned on Gen (18th Feb) 76%

Diesel Reading – 80%

Propane Reading – ~ 40%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 50%, ~ 276 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 205.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413342 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, to town.

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Screws tightened in the back airlock door. White tunnel covers were re-secured with new zipties, as well as tightening loose zipties and clipping the ends of the zipties.

Summary of Green Hab operations: Loft tank water filter replaced.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: We still need a 75W standard light bulb.